the area of a parallelogram and a square are the same. if the permeter of the square is 160 m and the heght of the parallalogram is 20 m , find the length of the corresponding base of the parallogram. 

Grade:7

2 Answers

Nishant Vora IIT Patna
7 years ago
side of square = 40 m

heght of the parallalogram is 20 m

hence area of square = 40*40 = 1600 m^2

so 1600 = base * 20

hence base = 80 m

Thanks
Kulsum Sheikh
7 years ago
The perimeter of square = 160 and Perimeter of square = 4 × side Therefore 160= 4 × side 160÷4 = side of square Side = 40 And so area of square will be (side)^2 Area of square = 40^2And we know that area of parallelogram = area of squareAnd area of parallelogram = base × height And height of parallelogram is 20 1600= base×201600÷20 = base of parallelogram Base = 80 cm Thankyou